OpenStudy (anonymous):

Where is the vertex of y = |3x - 6| + 3?

OpenStudy (anonymous):

Since we have an absolute value equation, then we have 2 equations.

OpenStudy (anonymous):

y1 = 3x - 6 + 3 = 3x -3 and y2 = -3x + 6 + 3 = -3x + 9

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

The vertex occurs when y1 = y2. That is 3x - 3 = -3x +9 6x = 12 x = 12/6 = 2 Replace x by 2 in y = |3x - 6| + 3, we get y = 3 Hence, the vertex is located at (2, 3)

OpenStudy (anonymous):

do you now?

OpenStudy (anonymous):

lol yea..just slow sometimes

OpenStudy (anonymous):

oh cool im glad i could help u bro